OpenStudy (xxstudyingstuffxx):

Solve for m. 12 + (7 +m) = 30 m = 11, correct?

OpenStudy (serenityjustice74):

12 + 7 + m = 30 19 + m = 30 m = 30 -19 = 11 SO YEAPPP ^=^

OpenStudy (xxstudyingstuffxx):

Thanks! ♥☺☻☺☻☺☻♥

OpenStudy (serenityjustice74):

No probes darling! ^-^

OpenStudy (xxstudyingstuffxx):

Darling? o.o